If a business borrows $16,000 and repays $23,600 in 5 years, what is the simple interest rate?
%

Answer by jim_thompson5910(35256)

A = P+Prt

23600 = 16000+16000r(5)

23600 = 16000+80000r

23600 - 16000 = 80000r

7600 = 80000r

7600/80000 = r

0.095 = r

r = 0.095


So the simple interest rate is 9.5%